OpenStudy (anonymous):

f(x)=2x+3^2 h(x)=3 what does (f(h(x))) equal what does (h(f(x))) equal

OpenStudy (anonymous):

f(h(x)) = 2(3)+3^2 = 6+3^2 h(f(x)) = NIL

OpenStudy (abb0t):

what does NIL mean?

OpenStudy (anonymous):

means nothing

OpenStudy (anonymous):

f(h(x))=f(x,x=h(x)=3)=6+9=15; h(f(x))=h(x,x=f(x))=3 since h(x) is a constant.